OpenStudy (anonymous):

guys help me! find the equation of the tangent line at (2,-1) on the circle (x+2)^2+(y+2)^2=25

OpenStudy (anonymous):

we know that the centre of the circle is at (-2,-2) from equation of circle. So the equation of the line (radius) joining (-2,-2) & (2,-1) y+2/x+2=-1/-4

OpenStudy (anonymous):

so y=x/4-3/2

OpenStudy (anonymous):

so slope of radius=1/4 the slope of the tangent which is perpendicular to the radius will be -4 is y=-4x+c. now since it passes through (2,-1)----> -1=-4*2+c

OpenStudy (bradely):

(x+2)^2+(y+2)^2=25 take derivative 2(x+2)+2(y+2)dy/dx =0 dy/dx=-(x+2)/(y+2) Slope at (2,-1) m= dy/dx=-4 y=mx+c y=-4x+c -1=-4(2)+c c=7 y=-4x+7 4x+y=7 Source: http://www.mathskey.com/question2answer/